OpenStudy (anonymous):

I need someone to explain this lesson to me. Be patient while I write the problem.

OpenStudy (anonymous):

Ok so it is walking me through the steps to solve for x \[\sqrt{x+21}-3=\sqrt{x}\]

OpenStudy (anonymous):

Here are the steps:\[\sqrt{x+21}=3+\sqrt{x}\] \[x+21=9+6\sqrt{x}+x\] \[12=6\sqrt{x}\] \[\sqrt{x}=2\] \[x=4\]

OpenStudy (anonymous):

The 2nd step is the one I didn't understand?\[x+21=9+6\sqrt{x}+x\]

OpenStudy (usukidoll):

ahhhh... they squared both sides.

OpenStudy (usukidoll):

because we want x... no x^2 stuff square root of x just x so we square both sides \[(\sqrt{x+21})^2=(3+\sqrt{x})^2\]

OpenStudy (anonymous):

Ok but wouldn't that leave us with \[x+21=9+x\]

OpenStudy (usukidoll):

the left side is correct.. the right one requires expansion

OpenStudy (usukidoll):

\[(\sqrt{x+21})(\sqrt{x+21}) = (3+\sqrt{x})(3+\sqrt{x})\]

OpenStudy (usukidoll):

so square root x square root is just whatever is in the square root so x+21 = (use foil method )

OpenStudy (anonymous):

Ohhh Ok I get it thank you because the OI of FOIL add together to get the \[6\sqrt{x}\] thank you

OpenStudy (usukidoll):

yup

OpenStudy (usukidoll):

then what they did was shift that x to the left to eliminate it and the 9 is shifted to the left

OpenStudy (usukidoll):

then divide both sides by 6 afterwards square both sides again and x = 4

OpenStudy (anonymous):

Yeah I get it thanks I am really liking this site it's helpin me out a lot